# Resolvent-Based Optimisation for Invariant Flows

### v1

@@ 5 Exact Solutions Found for RPCF
In this section the focus is a demonstration of the projected variational optimisation methodology applied to finding exact nonlinear solutions to the Navier-Stokes equations for RPCF. These solutions are obtained by providing a sufficiently large number of resolvent modes to the optimiser, such that basis can be considered complete. With this complete basis the global residual as defined in (6) can be made very small.
@@ 5.1 Implementation Details
This section discusses the numerics and programmatic strategies used to implement the procedure depicted in figure 3. All of the numerics are implemented in the Julia programming language with the majority of the code available at The-ReSolver. A single spatially and temporally extended scalar field is represented as a 3D array, with each dimension representing the wall-normal, spanwise, and time directions, respectively. The number of wall-normal points, spanwise points, and number of temporal samples are denoted by $N_{y}$, $N_{z}$, and $N_{t}$, respectively. Since the physical field is represented a real FFT transformed array the actual array has dimension of $N_{y}\times(\lfloor N_{z}/2\rfloor+1)\times N_{t}$, where $\lfloor\cdot\rfloor$ denotes the integer floor operation. The spanwise and time directions are transformed using FFTs (Frigo & Johnson, 2005), and so the discretisations in these directions are necessarily uniform. The distribution of wall-normal points is not restricted to be uniform since this direction is not represented with a Fourier transform. However, they are typically also uniform to allow easy interoperability with the DNS and Julia codes. The resolvent modes are represented by a 4D array, the extra dimension corresponding to multiple modes being available at each frequency and wall-normal location. The modal coefficients $a_{\bm{k}i}$ (and $s_{\bm{k}i}$) are represented very similarly in the program, only with the array dimension corresponding to the wall-normal direction now representing the modal index $i$, with a length $M$ for the number of modes used for the projection in (15) and (17). The storage of multiple fields with their complete time history, as well as the set of modes being used, represents the vast majority of the memory requirements of the programme. This problem is highly parallelisable due to the global representation, which is likely a necessary step to implementing these methods for fully 3D problems or for higher Reynolds numbers.
@@
Derivatives in the spanwise and time directions can be computed using standard spectral methods. Derivatives in the wall-normal direction are computed using finite difference methods, generally of fourth order accuracy. The finite difference stencils are derived using the custom package FDGrids.jl. To compute the projection in (15) and (17) it is necessary to compute integrals over all the directions given the definition of the inner-product in (3). In the spanwise and time directions this can be done by using Parseval’s theorem, converting the integral over these directions to sums of the coefficients over the corresponding wavenumbers. The wall-normal integrals are computed using the method of undetermined coefficients for quadratures (Dahlquist & Bj{̈o}rck, 2008). The nonlinear terms in the computations are computed using a pseudo-spectral method, using a $3/2$ padding rule in the spanwise and time directions to avoid aliasing errors.
@@
When constructing the modes $\bm{\psi}_{\bm{k}i}$ using resolvent analysis, the base flow used is the laminar solution $\bm{u}_{b}=y\hat{\bm{\imath}}$. This choice is made for its computational simplicity and numerical robustness; the laminar base flow is well-defined, analytically simple, and guarantees a tractable, well-conditioned singular value decomposition for all wavenumbers $\bm{k}\in\mathbb{Z}^{2}$ (including $\bm{k}=0$). Consequently, it provides a reliably calculable set of modes $\bm{\psi}_{\bm{k}i}$ without requiring a priori knowledge of the flow or introducing the numerical complications associated with a turbulent mean profile. For the computation of the SVD, the divide and conquer algorithm is used.
@@
We are free to choose the optimisation algorithm used for solving the optimisation problem (6), since any gradient-based method will work. Unless otherwise stated, L-BFGS is the algorithm used here. L-BFGS is a quasi-Newton method, incorporating approximate Hessian information into each iteration, which significantly improves the convergence properties of the optimisation near minima compared to gradient descent. This is particularly important here because the optimisation problem is non-convex, with many possible solutions potentially very close together. The problem is also generally quite poorly conditioned, mostly due to the influence of the Laplacian which has large effects on the high frequency components of the flow that decay away the most rapidly. L-BFGS’s ability to approximate the local curvature of the solution reduces the effect of this poor conditioning. Further discussions on each algorithm can be found in Nocedal & Wright (2006). The optimisation is performed using the Optim.jl package (Mogensen & Riseth, 2018). Any algorithm used here is also coupled with the Hager-Zhang line search algorithm as described in Hager & Zhang (2005). This further improves convergence rates, albeit at the possibility of reducing the robustness of the convergence. In practice, however, this is not observed to be an issue. The threshold for convergence is defined as $\mathcal{R}[\bm{u}]<10^{-12}$.
@@ 5.2 Equilibria
The Reynolds number to $\mathit{Re}=50$ and Rotation number $\mathit{Ro}=0.5$ for the equilibrium solutions. To begin, the optimiser is validated by observing its ability to reconstruct a known solution to the flow. This was obtained using the custom DNS solver, using a grid discretisation of $N_{y}=64$ on a uniform grid, and $N_{z}=32$ corresponding to $17$ spanwise Fourier modes. A set of $M=64$ resolvent modes $\bm{\psi}_{\bm{k}i}$ are used for this optimisation. To obtain the initial guess for the start of the optimisation S1 from DNS in figure 6 is projected onto the resolvent basis and then perturbed with random Gaussian white noise at each coefficient $a_{\bm{k}i}$. This results in the highly disordered field in figure 7(a). This perturbed flow is then optimised to try to recover the original solution. The result of this optimisation is shown in figure 7(b). This can be compared to S1 in figure 6, where the equilibria from the DNS and optimisation are qualitatively indistinguishable. Figure 7(c) shows the difference between the solution obtained from optimisation and from the DNS. The magnitude of the difference is a couple of orders magnitude smaller than that of the actual solution. The majority of the difference between the solutions is going to be a result of the differing numerics used in the DNS solver and Julia optimiser. This validates the optimiser possesses minima that correspond to solutions of (30).
@@
For comparison, the same initial condition was also optimised using the gradient descent and conjugate gradient algorithms (both still using Hager-Zhang line searches). The traces for each algorithm applied to the initial condition in figure 7(a) are shown in figure 8. Clearly L-BFGS performs the best, achieving the residual of $\mathcal{R}=10^{-12}$ after roughly $10000$ iterations. Gradient descent displays an initial large decrease in residual which quickly decreases as the convergence rate slows. This slow convergence rate is primarily due to the poor conditioning of the optimisation. Conjugate gradient also outperforms gradient descent, achieving a smaller residual for the same number of iterations. However, after the initial large decrease in the residual at the beginning of the optimisation, the convergence rate slows, approaching a similar speed as gradient descent. In addition, the iteration time of conjugate gradient is observed to be larger than L-BFGS.
@@
To investigate the robustness of the optimiser, it was initialised with various of initial conditions exciting certain spanwise wavenumbers with the intent of finding multiple new equilibrium solutions that are not observed in DNS of the flow at this Reynolds number. The same set of resolvent modes used for the previous optimisation are utilised here. The results of these optimisations are shown in figure 9. Figure 9(a) was initialised with the first $5$ modal coefficients $a_{\bm{k}i}$ excited with random values at the spanwise wavenumber of $k_{z}=4$, corresponding to streamwise rolls with half the wavelength of S1, with the rest of the coefficients left as zero. This converges to S3, one of the stable equilibrium shown in figure 6. Next the first $5$ modal coefficients are excited randomly for the spanwise wavenumber $k_{z}=1$, the solution in figure 9(b) is found, called S4(o). The addition of “(o)” in the name given to the equilibrium is in reference to the fact that it has been obtained from optimisation and is not a solution observed from the DNS. This solution again has the streamwise roll expected from the other equilibria solutions, but now with twice the spanwise wavelength. The highly symmetric structures of S3 and S4(o) closely mirror the solution of S1, and is ultimately expressing the same dynamics. To obtain a slightly more abnormal solution, the flow was initialised by exciting the first $8$ resolvent modes at the zeroth and first ($k_{z}=0,1$) spanwise wavenumbers with random values. The result of optimising such an initial condition is shown in figure 9(c), called S5(o). S5(o) shows a streamwise roll pattern with a width larger than in S1, roughly 3 spanwise units in length. This roll does not repeat over the remaining length of the domain since there is not sufficient room. Instead, the flow remains mostly stagnant in the remaining space, only expressing some very weak rolls that transports a small amount of the momentum from the top wall downwards.
@@
The spanwise power spectra of the intermediate flow over the duration of the optimisation leading to S3 are shown in figure 10. For clarity, the wavenumbers that are orders of magnitude smaller than those shown are filtered out. The filtered wavenumbers have negligible impact on the solution and are primarily a result of the spanwise domain size being large enough to contain multiple repetitions of the fundamental unit of the solution. The initial field used for the optimisation was excited at exactly one spanwise wavenumber, $k_{z}=4$. The optimiser initially increases the spectral power at the harmonic (and mean) wavenumbers as can be seen at iteration $i=20$, which is then followed by a gradual process of distributing the energy out from the higher wavenumbers until it settles on the final spectral peaks.
@@ 5.3 Periodic Solutions
The optimiser is now tasked with constructing a periodic solution at $\mathit{Re}=400$, a regime where RPCF displays a stable periodic motion as seen in figure 5. The flow is discretised with $N_{y}=128$, $N_{z}=64$, corresponding to $33$ spanwise modes, and $N_{t}=35$ temporal modes. A set of $M=64$ resolvent modes for each frequency are used for the optimisation. The smaller number of resolvent modes is used primarily to reduce computational effort. As will be seen in this section, the reduced number of modes compared to the degrees-of-freedom in the wall-normal direction have little effect on the accuracy of the final result. The initial guess for the optimiser is initialised in a similar way as in the first validation case in section 5.2. The periodic DNS data at $\mathit{Re}=400$ is projected onto the resolvent modes. The resulting coefficients $a_{\bm{k}i}$ are then perturbed with random Gaussian white noise. The resulting noisy flow is optimised.
Figure 11: Snapshots of the periodic solution ($\mathcal{R}\approx 5\times 10^{-13}$) obtained for $\mathit{Re}=400$ with a period of $T\approx 25.07$. Panel (a) shows $t=0$, panel (b) shows $t=T/4$, panel (c) shows $t=T/2$, and panel (d) shows $t=3T/4$.
Figure 11 shows a set of snapshots for the periodic solution obtained from optimisation at $\mathit{Re}=400$, sampled at points along its trajectory. The primary streamwise rolls are clearly present, and evolve in a wavy motion as consecutive vortices contract and expand. Figure 12 shows the spectrum of the periodic solution sampled at the two wall-normal positions $y\approx-0.86$ and $y\approx 0$. Each spectrum is a slice of the total spectrum such that only the positive temporal frequencies are shown, since the negative temporal frequencies are a reflection. An imbalance in the positive and negative temporal frequencies would manifest as travelling waves moving in the spanwise direction, which is not a feature of this solution. The general lack of spanwise motions in solutions to channel flows is a noted feature in Cvitanović & Gibson (2010). A chequerboard pattern is observed in the spectrum near the centreline of the flow, which represent the presence of the oscillating “tails” on either side of the vortices in figure 11. The most energetic mode is located at $\left(k_{z},\,k_{t}\right)=\left(1,\,0\right)$ which is mode that best fits the streamwise rolls that do not vary significantly with time. The spectrum at the location nearer the wall has a more continuous decay indicative of its less obvious spatial and temporal structure compared to the centre of the channel.
@@
Figure 13 shows the trace of the global residual and period in panels (a) and (b), respectively. The final residual achieved is roughly $\mathcal{R}\approx 3\times 10^{-13}$. The convergence rate of the problem is considerably slower after the initial rapid drop in residual than that observed in figure 8. This slow down can be explained by the increased degrees-of-freedom of the problem, both in terms of the inclusion of an extra temporal dimension and in terms of the added resolution required at higher Reynolds numbers. The DNS typically takes on the order of thousands of time units to decay to the periodic flow implying a rather small leading Floquet exponent. The potential for very weak linear stability has the effect of slowing convergence rates, as will be discussed in more detail in section 6.
@@ 6 Conditioning of the Optimisation
The convergence rate of variational optimisation algorithms slows considerably near a minimum, a problem governed by the local curvature of the residual function (Nocedal & Wright, 2006). The goal of this section is to link this conditioning directly to the underlying flow dynamics and to demonstrate that the resolvent modes provide a basis optimally constructed to improve the conditioning of the optimisation problem.
This curvature is characterized by the Hessian operator, which contains all second-derivative information at a point in optimisation space. The condition number of the Hessian measures the scale separation between the slowest and most rapid growth rates experienced by the objective when moving away from the minimum. A lower condition number implies improved convergence rates for gradient-based algorithms. A condition number of one signifies a perfectly quadratic local neighbourhood, while larger values indicate a long, narrow “valley”, which impedes convergence. It will be shown that the resolvent basis effectively minimizes this condition number.
@@
where $\bm{H}$ denotes the Hessian operator, $\mathcal{L}_{\bm{u}^{*}}$ denotes the linearised Navier-Stokes operator evaluated at the minimum $\bm{u}^{*}$, and $\mathbb{P}$ is the Leray projector as defined in section 3. The inclusion of the Leray projector is as a notational short-hand to avoid the added difficulty in having to explicitly account for the pressure gradient term in the Navier-Stokes equations. The detailed derivation of this equation is given in Appendix E. Let $\bm{v}$ be a unimodal perturbation to that minimum, i.e. let $\bm{v}=\bm{v}_{0}e^{i\omega t}$ for some arbitrary frequency where $\omega\in\mathbb{R}$ and some steady flow field $\bm{v}_{0}$. This perturbation can be substituted into (34), effectively performing a Fourier transform, which gives
@@
$0=\mu_{1}\leq\mu_{2}\leq\cdots\leq\mu_{n}\leq\cdots,$ (36)
and the condition number can be defined as $\kappa(\bm{H}_{\omega})=\mu_{N}/\mu_{1}$, where $\mu_{N}$ is the largest eigenvalue that can be resolved when the problem is discretised. The set of eigenvalues of $\bm{H}_{\omega}$ is unbounded since the eigenvalues of the Laplacian operator also unbounded.
It has been documented that the convergence rate of variational optimisers degrades near marginally stable solutions or bifurcation points (Farazmand, 2016; Lakoba & Yang, 2007). This phenomenon can be understood by considering the linear stability of the equilibrium solution $\bm{u}^{*}$, governed by the spectrum of $\mathcal{L}_{\bm{u}^{*}}$. Inspecting the derived expression for the Hessian in (35), it is clear that in the limit of any eigenvalue of $\mathcal{L}_{\bm{u}^{*}}$ approaching the imaginary axis (neutral stability), the action of the operator $i\omega\bm{I}-\mathcal{L}_{\bm{u}^{*}}$ becomes arbitrarily small near the frequency $\omega$ where that eigenvalue crosses. This implies that any perturbation $\bm{v}_{0}$ aligned with this marginally stable direction has a vanishingly small action from the Hessian, meaning the residual grows very slowly in this direction. In relation to the spectrum of $\bm{H}_{\omega}$, this is equivalent to the smallest Hessian eigenvalue $\mu_{1}$ approaching zero, which forces the condition number $\kappa(\bm{H}_{\omega})\to\infty$. This mathematical insight is the basis for acceleration methods that assume the optimiser is traversing along the most marginally stable mode (Azimi et al., 2022; Ashtari & Schneider, 2023), and several techniques have been devised to try and remove these problematic directions (Yang & Lakoba, 2007), although they have not been applied to fluid dynamics problems.
@@
Equation (38) represents an eigendecomposition of the Hessian, directly linking its spectrum to the resolvent of $\mathcal{L}{\bm{u}^{*}}$. The eigenvalues of $\bm{H}\omega$ are the squared inverse of the resolvent’s singular values ($\mu_{i}=\sigma_{\omega i}^{-2}$), and their corresponding eigenmodes are identical to the resolvent response modes.
Therefore, the condition number $\kappa(\bm{H}_{\omega})$ is primarily governed by the smallest retained singular value when the expansion is discretised. This relationship provides an optimal strategy for preconditioning since truncating the resolvent expansion to remove modes with the smallest singular values, one selectively eliminates the largest, most problematic eigenvalues from the Hessian. This optimally reduces the condition number, as it precisely targets and removes the directions associated with the strongest growth in the residual.
A key limitation is that these optimal resolvent modes depend on the linearised operator $\mathcal{L}_{\bm{u}^{*}}$ evaluated at the unknown minimum. Consequently, they cannot be used directly as the optimisation proceeds. The sensitivity of resolvent modes to the base flow remains poorly characterised, beyond the established dependence of singular values on the base flow (Brandt et al., 2011). We posit, however, that the higher-order modes exhibit universal structure, independent of the base flow, due to the dominance of viscous dissipation at small scales. This is favourable, as (38) indicates that rejecting these higher-order modes is the principal method for improving the Hessian’s condition number.

### v2

@@ 5 Exact Solutions Found for RPCF
In this section the focus is a demonstration of the projected variational optimisation methodology applied to finding exact nonlinear solutions to the Navier-Stokes equations for RPCF.
@@ 5.1 Implementation Details
This section discusses the numerics and programmatic strategies used to implement the procedure depicted in figure 3. All numerics are implemented in the Julia programming language, the majority of the code is available at The-ReSolver. A single spatially and temporally extended scalar field is represented as a 3D array, with each dimension representing the wall-normal, spanwise, and time directions, respectively. The number of wall-normal points, spanwise points, and temporal samples is denoted by $N_{y}$, $N_{z}$, and $N_{t}$, respectively. The actual array being stored for most computations, $\bm{u}_{\bm{k}}(y)$, has been transformed using a real FFT and thus has dimensions of $N_{y}\times(\lfloor N_{z}/2\rfloor+1)\times N_{t}$, where $\lfloor\cdot\rfloor$ denotes the integer floor operation. The spanwise and time directions are transformed using FFTW (Frigo & Johnson, 2005), and so the discretisations in these directions are necessarily uniform. In this work the wall-normal points are also uniformly distributed to allow for easy interoperability between the discretisations used by the DNS and Julia codes. The resolvent modes are represented by a 4D array, the extra dimension corresponding to the multiple modes available at each frequency and wall-normal location. This mode array therefore has dimensions of $3N_{y}\times M\times(\lfloor N_{z}/2\rfloor+1)\times N_{t}$ where $M\in\mathbb{N}$ is the number of modes used for the Galerkin projection. The modal coefficients $a_{\bm{k}i}$ (and $s_{\bm{k}i}$) are represented very similarly to the spectral arrays, only with the array dimension corresponding to the wall-normal direction now representing the modal index $i$, with a length $M$ for the number of modes used for the projection in (15) and (17). Thus, these arrays have a size of $M\times(\lfloor N_{z}/2\rfloor+1)\times N_{t}$. Since $M\sim N_{y}$ when the goal is to find exact solutions, the memory footprint of the coefficient arrays $a_{\bm{k}i}$ and $s_{\bm{k}i}$ is similar to the spectral arrays $\bm{u}_{\bm{k}}(y)$. The largest component of the memory requirements for finding exact solutions comes from the modes themselves, where all dimensions of the array grow with Reynolds number. This problem is highly parallelisable due to the global representation, which would be a necessary step to implementing these methods for fully 3D problems or for higher Reynolds numbers.
@@
Derivatives in the spanwise and time directions can be computed using standard spectral methods. Derivatives in the wall-normal direction are computed using finite difference methods, generally of second order accuracy to match as closely with the numerics of the DNS solver. The finite difference stencils are derived using the custom package FDGrids.jl. To compute the projection in (15) and (17) it is necessary to compute integrals over all the directions given the definition of the inner-product in (3). In the spanwise and time directions this can be done by using Parseval’s theorem, converting the integral over these directions to sums of the coefficients over the corresponding wavenumbers. The wall-normal integrals are computed using the method of undetermined coefficients for quadratures (Dahlquist & Björck, 2008). The nonlinear terms in the computations are computed using a pseudo-spectral method, using a $3/2$ padding rule in the spanwise and time directions to avoid aliasing errors.
@@
When constructing the modes $\bm{\psi}_{\bm{k}i}$ using resolvent analysis, the base flow used is the laminar solution $\bm{u}_{b}=y\hat{\bm{\imath}}$. This choice is made for its computational simplicity and numerical robustness; the laminar base flow is well-defined, analytically simple, and guarantees a tractable, well-conditioned singular value decomposition for all wavenumbers $\bm{k}\in\mathbb{Z}^{2}$ (including $\bm{k}=0$). Consequently, it provides a reliably calculable set of modes $\bm{\psi}_{\bm{k}i}$ without requiring a priori knowledge of the flow or introducing the numerical complications associated with a turbulent mean profile. It has been observed that the difference in the modes generated using the laminar profile and turbulent mean are not significantly different, thus the optimisation is not greatly affected by this choice. For the computation of the SVD, the divide and conquer algorithm is used.
@@
We are free to choose the optimisation algorithm used for solving the optimisation problem (6), since any gradient-based method will work. Unless otherwise stated, L-BFGS is the algorithm used here. L-BFGS is a quasi-Newton method, incorporating approximate Hessian information into each iteration, which significantly improves the convergence properties of the optimisation near minima compared to gradient descent. This is particularly important here because the optimisation problem is non-convex, with many possible solutions potentially very close together. The problem is also generally quite poorly conditioned, a topic that will be further expanded upon in section 6. L-BFGS’s ability to approximate the local curvature of the solution reduces the effect of this poor conditioning. Further discussions on each algorithm can be found in Nocedal & Wright (2006). The optimisation is performed using the Optim.jl package (Mogensen & Riseth, 2018). Any algorithm used here is also coupled with the Hager-Zhang line search algorithm as described in Hager & Zhang (2005). This further improves convergence rates, albeit at the possibility of reducing the robustness of the convergence. In practice, however, this is not observed to be an issue. The threshold for convergence is defined as $\mathcal{R}[\bm{u}]<10^{-12}$.
@@ 5.2 Equilibria
The Reynolds number is set to $\mathit{Re}=50$ and rotation number $\mathit{Ro}=0.5$ for the equilibrium solutions. To begin, the optimiser is validated by observing its ability to reconstruct a known solution to the flow. This was obtained using the custom DNS solver, with a grid discretisation of $N_{y}=64$ on a uniform grid, and $N_{z}=32$ corresponding to $17$ spanwise Fourier modes. A set of $M=64$ resolvent modes $\bm{\psi}_{\bm{k}i}$ are used for this optimisation. To obtain the initial guess for the start of the optimisation, S1 from DNS in figure 5 is projected onto the resolvent basis and then perturbed with random Gaussian white noise at each coefficient $a_{\bm{k}i}$. This results in the highly disordered field in figure 6(a). This perturbed flow is then optimised to try to recover the original solution. The result of this optimisation is shown in figure 6(b). This can be compared to S1 in figure 5, where the equilibria from the DNS and optimisation are qualitatively indistinguishable. Figure 6(c) shows the difference between the solution obtained from optimisation and from the DNS. The magnitude of the difference is a couple of orders magnitude smaller than that of the actual solution. The difference between the solutions is a result of the differing numerics used in the DNS solver and Julia optimiser. This validates the optimiser possesses minima that correspond to solutions of (30).
@@
For comparison, the same initial condition was also optimised using the gradient descent and conjugate gradient algorithms (both still using Hager-Zhang line searches). The traces for each algorithm applied to the initial condition in figure 6(a) are shown in figure 7. Clearly L-BFGS performs the best, achieving the residual of $\mathcal{R}=10^{-12}$ after a little over $4000$ iterations. Gradient descent displays an initial large decrease in residual which quickly decreases as the convergence rate slows. This final and slow phase of the optimisation is a result of the optimiser approaching the minimum along the direction associated with the slowest growth direction in $\mathcal{R}$. Conjugate gradient also outperforms gradient descent, achieving a smaller residual for the same number of iterations. However, after the initial large decrease in the residual at the beginning of the optimisation, the convergence rate slows, approaching a similar speed as gradient descent. In addition, the iteration time of conjugate gradient is observed to be larger than L-BFGS.
@@
To investigate the robustness of the optimiser, it was initialised with initial conditions exciting certain spanwise wavenumbers with the intent of finding multiple new equilibrium solutions that are not observed in DNS at this Reynolds number. The same set of resolvent modes used for the previous optimisation are utilised here. The results of these optimisations are shown in figure 8. Figure 8(a) was initialised with the first 33rd to $5$th modal coefficients $a_{\bm{k}i}$ excited with random values at the spanwise wavenumber of $k_{z}=4$, corresponding to streamwise rolls with half the wavelength of S1, along with as the 33rd mode at $k_{z}=0$, with the rest of the coefficients left as zero. This synthetically constructed initial flow converges to S3, one of the stable equilibria shown in figure 5. Next, when the first $5$ modal coefficients are excited randomly for the spanwise wavenumber $k_{z}=1$, the solution in figure 8(b) is found, called S4(o). The addition of “(o)” in the name given to the equilibrium is in reference to the fact that it has been obtained from optimisation and is not a solution observed from the DNS, a result of them likely being linearly unstable. This solution again has the streamwise roll expected from the other equilibria solutions, but now with twice the spanwise wavelength. The highly symmetric structures of S3 and S4(o) closely mirror the solution of S1, and is ultimately expressing the same dynamics. To obtain a slightly more abnormal solution, the flow was initialised by exciting the first $8$ resolvent modes at the zeroth and first ($k_{z}=0,1$) spanwise wavenumbers with random values. The result of optimising such an initial condition is shown in figure 8(c), called S5(o). S5(o) shows a streamwise roll pattern with a width larger than in S1, roughly 3 spanwise units in length. This roll does not repeat over the remaining length of the domain since there is not sufficient room. Instead, the flow remains mostly stagnant in the remaining space, only expressing some very weak rolls that transports a small amount of the momentum from the top wall downwards.
@@
The spanwise power spectra of the intermediate flow over the duration of the optimisation leading to S3 are shown in figure 9. For clarity, the wavenumbers that are orders of magnitude smaller than those shown are omitted. The omitted wavenumbers have negligible impact on the solution and are primarily a result of the spanwise domain size being large enough to contain multiple repetitions of the fundamental unit of the solution. The initial field used for the optimisation was excited at exactly one spanwise wavenumber, $k_{z}=4$. The optimiser initially spreads out the spectral content from the initial condition to the other wavenumbers as can be seen at iteration $i=20$. The large wavenumbers are then damped by viscous effects and the energy containing low wavenumbers are tuned towards the final solution by $i=500$. The remaining iterations are focused on finely tuning the higher wavenumbers while gradually decreasing the residual, until a balanced flow is achieved at $i=2200$.
@@ 5.3 Periodic Solutions
The optimiser is now tasked with constructing a periodic solution at $\mathit{Re}=400$, a regime where RPCF displays a stable periodic motion as seen in figure 4. The flow is discretised with $N_{y}=128$, $N_{z}=64$, corresponding to $33$ spanwise modes, and $N_{t}=35$ temporal modes. A set of $M=64$ resolvent modes for each frequency are used for the optimisation. The smaller number of resolvent modes is used primarily to reduce computational effort. As will be seen in this section, the reduced number of modes compared to the degrees-of-freedom in the wall-normal direction has little effect on the accuracy of the final result. The initial guess for the optimiser is initialised in a similar way as in the first validation case in section 5.2. The periodic DNS data at $\mathit{Re}=400$ is projected onto the resolvent modes. The resulting coefficients $a_{\bm{k}i}$ are then perturbed with random Gaussian white noise and the resulting noisy flow is optimised.
Figure 10: Snapshots of the periodic solution ($\mathcal{R}\approx 5\times 10^{-11}$) obtained for $\mathit{Re}=400$ with a period of $T\approx 25.05$. Panel (a) shows $t=0$, panel (b) shows $t=T/4$, panel (c) shows $t=T/2$, and panel (d) shows $t=3T/4$.
Figure 10 shows a set of snapshots for the periodic solution obtained from optimisation at $\mathit{Re}=400$, sampled at points along its trajectory. The primary streamwise rolls are clearly present, and evolve in a wavy motion as consecutive vortices contract and expand. Figure 11 shows the spectrum of the periodic solution sampled at the two wall-normal positions $y\approx-0.86$ and $y\approx 0$. Each spectrum is a slice of the total spectrum such that only the positive temporal frequencies are shown, since the negative temporal frequencies are a reflection. An imbalance in the positive and negative temporal frequencies would manifest as travelling waves moving in the spanwise direction, which is not a feature of this solution. The general lack of spanwise motions in solutions to channel flows is a noted feature in Cvitanović & Gibson (2010). A chequerboard pattern is observed in the spectrum near the centreline of the flow, which represent the presence of the oscillating “tails” on either side of the vortices in figure 10. The most energetic mode is located at $\left(k_{z},\,k_{t}\right)=\left(1,\,0\right)$ which is the mode that best fits the streamwise rolls that do not vary significantly with time. The spectrum at the location nearer the wall has a more continuous decay indicative of its less obvious spatial and temporal structure compared to the centre of the channel.
@@
Figure 12 shows the trace of the global residual and period in panels (a) and (b), respectively. The final residual achieved is roughly $\mathcal{R}\approx 5\times 10^{-11}$. The convergence rate of the problem is considerably slower after the initial rapid drop in residual than that observed in figure 7. This slower convergence compared to that observed in figure 7 can be explained by the increased degrees-of-freedom of the problem, both in terms of the inclusion of an extra temporal dimension and in terms of the added resolution required at higher Reynolds numbers. As a final point, the maximal and minimal kinetic energies of the optimised periodic solution is plotted on figure 4 (denoted with red triangles) with the corresponding kinetic energy extrema from DNS, which shows good agreement between the solutions from the two different numerical sources.
@@ 6 Conditioning of the Optimisation
In the literature the convergence rate of variational optimisation algorithms slows considerably near a minimum, a problem governed by the local curvature of the residual function (Nocedal & Wright, 2006). This curvature is characterized by the Hessian operator, which contains all second-derivative information at a point in optimisation space. The condition number of the Hessian measures the scale separation between the slowest and most rapid growth rates experienced by the objective when moving away from the minimum. A lower condition number implies improved convergence rates for gradient-based optimisation algorithms. A condition number of one signifies a perfectly quadratic local neighbourhood, while larger values indicate a long, narrow “valley”, which impedes convergence. The goal of this section is to link this conditioning directly to the underlying flow dynamics and to demonstrate that the resolvent modes provide a basis optimally constructed to improve the conditioning of the optimisation problem.
@@
where $\bm{H}$ denotes the Hessian operator, $\mathcal{L}_{\bm{u}^{*}}$ denotes the linearised Navier-Stokes operator evaluated at the minimum $\bm{u}^{*}$, and $\mathbb{P}$ is the Leray projector as defined in section 3. The inclusion of the Leray projector is as a notational short-hand to avoid the added difficulty in having to explicitly account for the pressure gradient term in the Navier-Stokes equations. The detailed derivation of this equation is given in Appendix E. Let $\bm{v}$ be a unimodal perturbation to that minimum, i.e. let $\bm{v}=\bm{v}_{0}e^{i\omega t}+\text{c.c.}$ for some arbitrary frequency where $\omega\in\mathbb{R}$ and some steady flow field $\bm{v}_{0}$, where c.c. denotes the complex conjugate. This perturbation can be substituted into (34), effectively performing a Fourier transform, which gives
@@
$0\leq\mu_{\omega 1}\leq\mu_{\omega 2}\leq\cdots\leq\mu_{\omega n}\leq\cdots,$ (36)
and the condition number can be defined as $\kappa(\bm{H}_{\omega})=\mu_{\omega N}/\mu_{\omega 1}$, where $\mu_{\omega N}$ is the largest eigenvalue that can be resolved when the problem is discretised. The set of eigenvalues of $\bm{H}_{\omega}$ is unbounded since the eigenvalues of the Laplacian operator are also unbounded.
It has been documented that the convergence rate of variational optimisers degrades near marginally stable solutions or bifurcation points (Farazmand, 2016; Lakoba & Yang, 2007). This phenomenon can be understood by considering the linear stability of the equilibrium solution $\bm{u}^{*}$, governed by the spectrum of $\mathcal{L}_{\bm{u}^{*}}$. Inspecting the derived expression for the Hessian in (35), it is clear that in the limit of any eigenvalue of $\mathcal{L}_{\bm{u}^{*}}$ approaching the imaginary axis (neutral stability), the action of the operator $i\omega\bm{I}-\mathcal{L}_{\bm{u}^{*}}$ becomes arbitrarily small near the frequency $\omega$ where that eigenvalue crosses. This implies that any perturbation $\bm{v}_{0}$ aligned with this marginally stable direction has a vanishingly small action from the Hessian, meaning the residual grows very slowly in this direction. In relation to the spectrum of $\bm{H}_{\omega}$, this is equivalent to the smallest Hessian eigenvalue $\mu_{\omega 1}$ approaching zero, which forces the condition number $\kappa(\bm{H}_{\omega})\to\infty$. This mathematical insight is the basis for acceleration methods that assume the optimiser is traversing along the most marginally stable mode (Azimi et al., 2022; Ashtari & Schneider, 2023), and several techniques have been devised to try and remove these problematic directions (Yang & Lakoba, 2007), although they have not been applied to fluid dynamics problems.
@@
Equation (38) represents an eigendecomposition of the Hessian, directly linking its spectrum to the resolvent of $\mathcal{L}_{\bm{u}^{*}}$. The eigenvalues of $\bm{H}_{\omega}$ are the squared inverse of the resolvent’s singular values ($\mu_{\omega i}=\sigma_{\omega i}^{-2}$), and their corresponding eigenmodes are identical to the resolvent response modes. Therefore, the condition number $\kappa(\bm{H}_{\omega})$ is primarily governed by the smallest retained singular value when the expansion is discretised. This relationship provides an optimal strategy for preconditioning since by truncating the resolvent expansion to remove the modes with the smallest singular values, one selectively eliminates the largest, most problematic eigenvalues from the Hessian. This optimally reduces the condition number, as it precisely targets and removes the directions associated with the strongest growth in the residual. In Mons & Marquet (2021) the authors attempted to derive an methodology to determine optimal sensor placement for data assimilation. Their formulation lead to a similar relationship to (37) relating the Hessian operator of their problem to the resolvent of their linearised dynamics. The fact that this relationship appears in both this work and Mons & Marquet (2021), despite the different objectives, is indicative of the role the resolvent has in modelling the higher order sensitivities in these types of optimisation problems.
A key limitation is that these optimal resolvent modes depend on the linearised operator $\mathcal{L}_{\bm{u}^{*}}$ evaluated at the unknown minimum. Thus, to be able to derive the modes used in (38) the solution has to be known before the optimisation. Therefore it appears that this does not provide a useful method for improving convergence rates. We posit, however, that the higher-order modes exhibit universal structure, independent of the base flow, due to the dominance of viscous dissipation at small scales. This is favourable, as (38) indicates that rejecting these higher-order modes is the principal method for improving the Hessian’s condition number. It should be noted that the sensitivity of resolvent modes to the base flow used for the linearisation remains poorly characterised, beyond the established dependence of the singular values in Brandt et al. (2011).